A firm's demand curve is P = 1 - 2Q. The firm has a current price of R1000 and it sells 100 units per day. What is the firm's price elasticity of demand?

1
Expert's answer
2022-04-19T15:23:37-0400

P = 1 - 2Q, so Q = 0.5 - 0.5P.

The firms price elasticity of demand is:

"Ed=\u22120.5\u00d71000\/100=\u22125."

So, the demand is elastic at this price.